1. Find the coordinates of the point which divides the join of (- 1, 7) and (4, – 3) in the ratio 2:3.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let P(x, y) be the required point. Using the section formula, we get x = (2×4 + 3×(-1))/(2 + 3) = (8 – 3)/5 = 1 y = (2×-3 + 3×7)/(2 + 3) = (-6 + 21)/5 = 3 Therefore, the point is (1, 3).